The length of a rectangle is 7 inches longer than it is wide. If the area is 60 square inches, what are the dimensions of the rectangle?

The width, or shorter side is ___ inches

The length, or longer side is ___ inches
Answer #1

I 7th det width So length 7+ W Area = so square? = 60 -> (1N) (7+) NA TW-60=0 w²t =) 2 W +124-50-60=0 w (i+12)-5 (1-1+12 ) =